OpenStudy (albert0898):

An object detected on radar is 5 miles to the east, 4 miles to the north, and 1 mile above the tracking station. Among the following, which is the closest approximation to the distance, in miles, that the object is from the tracking station? A. 6.5 B. 7.2 C. 8.3 D. 9.0 E. 10.0 Please explain the steps. You'll be rewarded!

OpenStudy (taramgrant0543664):

Do you know how to figure this out?

OpenStudy (taramgrant0543664):

Your first step is to find the hypotenuse on the triangle using the distance east as one side and the distance north as the other side. I will use c to represent the length of the hypotenuse. c^2=(5^2)(4^2) Can you solve for c?

OpenStudy (taramgrant0543664):

c^2=(25)(16) c^2=400 c=sqrt(400) c=?

OpenStudy (taramgrant0543664):

I can walk you through this @Albert0898 but I can't just give you the answer you have to do some work to here

OpenStudy (albert0898):

Sorry, I was afk. c is 20

OpenStudy (taramgrant0543664):

That is a okay and yes c is 20 so now we can make a new triangle the one side being 20 and the other will be the distance up which is 1mile so we can do the same as before by finding the hypotenuse lets call this one d d^2=(20^2)(1^2) d^2=(400)(1) d=?

OpenStudy (albert0898):

d = 20

OpenStudy (taramgrant0543664):

Yes that would be right

OpenStudy (taramgrant0543664):

Wait a second I screwed up the formula c^2=(5^2)+(4^2) c^2=25+16 c^2=41 c=6.4 so d^2=(6.4^2)+(1^2) d=6.4 I'm sorry I really don't know how I managed to forget the plus sign!!!!!!!!!!